In a 24-hour endurance race, a car makes 640 laps of a race track that is 6 km per lap. How far did the car travel in 24 hours? What was its average speed?

1
Expert's answer
2022-03-29T12:28:04-0400

The distance traveled by a car

"d=Nd_1\\
=640\times 6\rm\: km=3840\:\rm km"

The average speed

"v=\frac{d}{t}\\
=\frac{3840\:\rm km}{24\:\rm hr}=160\:\rm km/hr"